OpenStudy (anonymous):

Suppose that y varies inversely with x. Write an equation for the inverse variation. y = 2 when x = 5

OpenStudy (solomonzelman):

\(\Large\color{red}{ \bf y~~varies~~inversely~~with~~x~~means, }\) \(\Large\color{blue}{ \bf y= \frac{k}{x} }\) where x is the constant of variation.

OpenStudy (solomonzelman):

saying that \(\Large\color{blue}{ \bf 2= \frac{k}{5} }\) that gives k=10

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf y= \frac{10}{x} }\) would be your equation for any x.

OpenStudy (anonymous):

i was just about to say that thanks again i have another one like that if i slove it can you just tell me if im right or wrong please?

OpenStudy (anonymous):

y = 2.5 when x = 9 so its k = 22.5; y = 22.5x?

OpenStudy (solomonzelman):

direct or inverse variation ?

OpenStudy (anonymous):

inverse variation.

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf y= \frac{k}{x} }\) \(\Large\color{blue}{ \bf 2.5= \frac{k}{9} }\) this gives k=22.5 thus you get \(\Large\color{blue}{ \bf y= \frac{22.5}{x} }\)

OpenStudy (anonymous):

so im right?

OpenStudy (anonymous):

Find the constant of variation k for the inverse variation. Then write an equation for the inverse variation. y = 2.5 when x = 9

OpenStudy (anonymous):

the answer is what i just said right?